1.  A sample of gas has a pressure of 1.45 atm at a temperature of 22 C. What    would the pressure of the gas be if it is heated to 37 C?




1
Expert's answer
2021-09-30T03:24:57-0400

P1/T1= P2/T2

1.45/295=X/310

0.00492= x/310

X = 0.00492×310

= 1.52atm